# Efficient Spin-Adapted Implementation of Multireference Algebraic Diagrammatic Construction Theory. I. Core-Ionized States and X-Ray Photoelectron Spectra

### v1

@@ 5.2 Azobenzene Photoisomerization
We performed the MR-ADC(2)-X calculations of XPS spectra at three azobenzene geometries: the equilibrium structures of trans- and cis-isomers and the geometry of minimum energy conical intersection (MECI) between the two lowest-energy singlet states that is believed to be important in photoisomerization (Section 4).218, 204, 213 For each structure, the carbon and nitrogen K-edge XPS signatures were simulated for the ground () and lowest-energy excited () singlet states. For the trans- and cis-isomers, the state corresponds to the excitation from the molecular orbital localized on nitrogen lone pairs to the lowest-energy -antibonding orbital of the molecule (Figures 4b, S6, S9). We also computed the XPS spectra of trans-azobenzene in the second singlet excited state () corresponding to the electronic transition (Figures 4c and S7). The calculations employed large quadruple-zeta basis sets correlating all electrons in 1476 molecular orbitals and CAS(16e,15o). Additional computational details can be found in Section 4 and Supplementary Information. We note that our calculations did not incorporate excited-state relaxation and nuclear dynamics effects that can be important for accurate interpretation of TR-XPS spectra.
Figures 3a to 3c show the simulated C K-edge XPS spectra of trans- and cis-azobenzene in the , , and electronic states. For the ground state of trans-isomer, the results of simulations are in an excellent agreement with the experimental spectrum (Figure 3d) measured by Carlini et al.215 In all C K-edge XPS spectra, an intense peak is followed by a feature with weaker intensity that appears as a shoulder in the spectrum. These two features correspond to the K-edge ionization of C atoms in two distinct chemical environments: 1) ten H-bonded C’s and 2) two N-bonded C’s. Due to increasing electronegativity () in the order (H) (C) (N), the N-bonded carbons exhibit a higher charge and less efficient core-hole screening than those attached to the H atoms, resulting in a significant blueshift of the corresponding feature. As can be seen in Figure 3c, the ground-state XPS spectra of trans- and cis-isomers exhibit nearly identical spacing between the two peaks ( 0.6 eV), indicating that the electronic environment of C atoms in these states is similar.
@@
Figures 3e to 3g present the N K-edge XPS spectra of trans- and cis-azobenzene. For each structure and electronic state, the N K-edge spectrum displays a single peak corresponding to the 1s ionization of two symmetry-equivalent nitrogen atoms. The MR-ADC(2)-X method underestimates the experimental core binding energy215 of trans-azobenzene in the state by 0.6 eV. In contrast to the C K-edge XPS spectra, the computed N K-ionization energies are more sensitive to the cis/trans-orientation of ground state geometry or excitation to the state. Specifically, isomerization from trans- to cis-azobenzene on the potential energy surface increases the N K-binding energy by 0.5 eV while the excitation at the trans geometry decreases it by 0.8 eV. Exciting the molecule to the state results in 0.2 and 0.4 eV blueshifts in peak position for the trans- and cis-isomers, respectively. As for the C K-edge, all spectral changes can be interpreted based on the analysis of CASSCF natural orbitals (Figure 3). For example, the significant ( 0.8 eV) blueshift of N K-signal for the state is consistent with excited-state electron density concentrating on the orbitals localized on the N atoms ( and , Figure 3c) as opposed to the orbitals localized on the phenyl rings ( and ).
@@
Finally, to demonstrate an application of MR-ADC at non-equilibrium regions of potential energy surfaces, we analyze the XPS spectra computed for the – MECI geometry shown in Figure 5a. The MECI exhibits an unusual structure of symmetry with a 92.7 dihedral angle between the two phenyl rings (Ph) and significantly different N–C bond distances (1.40 and 1.33 Å). As can be seen from the occupations of CASSCF natural orbitals (Figures S10 and S11), the two singlet states participating in MECI show significantly different degree of open-shell character. Consistent with Ref. 217, we will refer to the state with lower/higher open-shell character as “closed-shell”/“open-shell” and label them as /. To distinguish between the two non-equivalent NPh fragments, we denote the Ph group with the shorter/longer N–C bond as PhD/PhA. The Mulliken analysis of charge density (Figure 5b,c) reveals that NPhD has a lower total charge than NPhA in the state, but a higher charge in the state.
@@
The C and N K-edge XPS spectra of and simulated at the MECI geometry are shown in Figures 5d to 5g. Due to the asymmetric structure, the spectra show distinct XPS signatures for the C and N atoms of the NPhD and NPhA groups of the molecule. In the closed-shell state, the lowest-energy C and N K-edge ionization occurs in the PhD fragment, which exhibits significantly higher electron density and more efficient core-hole screening. Ionizing the PhA group requires more than 1 eV of additional ionization energy, significantly broadening the C and N K-edge XPS spectra in comparison to those of ground-state azobenzene (Figure 3). The XPS signatures of open-shell state resemble the ground-state spectra much more closely with overlapping peaks originating from PhD and PhA. These results are consistent with the analysis of Mulliken charges (Figures 5b and 5c) and CASSCF natural orbitals (Figures S10 and S11), which reveal a more even electron density distribution between NPhD and NPhA in the state.
@@ 6 Conclusions
In this work, we presented an efficient implementation of multireference algebraic diagrammatic construction theory with core-valence separation for simulating core-ionized states and X-ray photoelectron spectra (CVS-IP-MR-ADC). Developed in the open-source and freely available Prism program,86 CVS-IP-MR-ADC takes advantage of spin adaptation, automatic code generation, and efficient handing of two-electron integrals via density fitting. Incorporating dynamic and static correlation, CVS-IP-MR-ADC allows to accurately simulate the X-ray or extreme ultraviolet photoelectron spectra (XPS) for molecules in excited electronic states or at nonequilibrium ground-state geometries. The CVS-IP-MR-ADC is resilient to intruder-state problems and allows to calculate many (10’s or even 100’s) excited states starting with a single CASSCF wavefunction for the reference electronic state.
We demonstrated the capabilities of our efficient CVS-IP-MR-ADC implementation by applying it to substituted ferrocene complexes and azobenzene molecule along its photoisomerization pathway. In all calculations, we used core-polarized quadruple-zeta basis sets, correlating all electrons in more than 1500 molecular orbitals. For the ground electronic states of substituted ferrocenes and trans-azobenzene, the XPS spectra simulated using the extended second-order CVS-IP-MR-ADC method (CVS-IP-MR-ADC(2)-X) are in excellent agreement with experimental measurements. Encouraged by these results, we also computed the XPS spectra for the azobenzene molecule in its excited states and at the geometry of asymmetric minimum energy conical intersection, which may be useful in interpreting the future time-resolved XPS experiments.
The efficient implementation strategy presented in this work can be used to develop fast MR-ADC methods for simulating other spectroscopic properties, including electron attachment and ionization in the UV/Vis region,37, 38 and neutral excitations in the UV/Vis and X-ray absorption spectroscopies.24, 35, 36 These developments will be reported in the forthcoming publications of this series and will further expand the application domain of MR-ADC.
This work was supported by the National Science Foundation, under Grant No. CHE-2044648. The authors would like to thank Nicholas Gaba for testing the CVS-IP-MR-ADC code and valuable suggestions on its improvements. Computations were performed at the Ohio Supercomputer Center under Project No. PAS1583.219

### v2

@@ 5.2 Azobenzene Photoisomerization
We performed the MR-ADC(2)-X calculations of XPS spectra at three azobenzene geometries: the equilibrium structures of trans- and cis-isomers and the geometry of minimum energy conical intersection (MECI) between the two lowest-energy singlet states that is believed to be important in photoisomerization (Section 4).224, 210, 219 For each structure, the carbon and nitrogen K-edge XPS signatures were simulated for the ground () and lowest-energy excited () singlet states. For the trans- and cis-isomers, the state corresponds to the excitation from the molecular orbital localized on nitrogen lone pairs to the lowest-energy -antibonding orbital of the molecule (Figures 4b, S6, S9). We also computed the XPS spectra of trans-azobenzene in the second singlet excited state () corresponding to the electronic transition (Figures 4c and S7). The calculations employed large quadruple-zeta basis sets correlating all electrons in 1476 molecular orbitals and CAS(16e,15o). Additional computational details can be found in Section 4 and Supplementary Information. We note that our calculations did not incorporate excited-state relaxation and nuclear dynamics effects that can be important for accurate interpretation of TR-XPS spectra.
Figures 3a to 3c show the simulated C K-edge XPS spectra of trans- and cis-azobenzene in the , , and electronic states. For the ground state of trans-isomer, the results of simulations are in an excellent agreement with the experimental spectrum (Figure 3d) measured by Carlini et al.221 In all C K-edge XPS spectra, an intense peak is followed by a feature with weaker intensity that appears as a shoulder in the spectrum. These two features correspond to the K-edge ionization of C atoms in two distinct chemical environments: 1) ten H-bonded C’s and 2) two N-bonded C’s. Due to increasing electronegativity () in the order (H) (C) (N), the N-bonded carbons exhibit a higher charge and less efficient core-hole screening than those attached to the H atoms, resulting in a significant blueshift of the corresponding feature. As can be seen in Figure 3c, the ground-state XPS spectra of trans- and cis-isomers exhibit nearly identical spacing between the two peaks ( 0.6 eV), indicating that the electronic environment of C atoms in these states is similar.
@@
Figures 3e to 3g present the N K-edge XPS spectra of trans- and cis-azobenzene. For each structure and electronic state, the N K-edge spectrum displays a single peak corresponding to the 1s ionization of two symmetry-equivalent nitrogen atoms. The MR-ADC(2)-X method underestimates the experimental core binding energy221 of trans-azobenzene in the state by 0.6 eV. In contrast to the C K-edge XPS spectra, the computed N K-ionization energies are more sensitive to the cis/trans-orientation of ground state geometry or excitation to the state. Specifically, isomerization from trans- to cis-azobenzene on the potential energy surface increases the N K-binding energy by 0.5 eV while the excitation at the trans geometry decreases it by 0.8 eV. Exciting the molecule to the state results in 0.2 and 0.4 eV blueshifts in peak position for the trans- and cis-isomers, respectively. As for the C K-edge, all spectral changes can be interpreted based on the analysis of CASSCF natural orbitals (Figure 3). For example, the significant ( 0.8 eV) blueshift of N K-signal for the state is consistent with excited-state electron density concentrating on the orbitals localized on the N atoms ( and , Figure 3c) as opposed to the orbitals localized on the phenyl rings ( and ).
@@
Finally, to demonstrate an application of MR-ADC at non-equilibrium regions of potential energy surfaces, we analyze the XPS spectra computed for the – MECI geometry shown in Figure 5a. The MECI exhibits an unusual structure of symmetry with a 91.6 dihedral angle between the two phenyl rings (Ph) and significantly different N–C bond distances (1.39 and 1.35 Å). As can be seen from the occupations of CASSCF natural orbitals (Figures S10 and S11), the two singlet states participating in MECI show significantly different degree of open-shell character. Consistent with Ref. 223, we will refer to the state with lower/higher open-shell character as “closed-shell”/“open-shell” and label them as /. To distinguish between the two non-equivalent NPh fragments, we denote the Ph group with the shorter/longer N–C bond as PhD/PhA. The Mulliken analysis of charge density (Figure 5b,c) reveals that NPhD has a lower total charge than NPhA in the state, but a higher charge in the state.
@@
The C and N K-edge XPS spectra of and simulated at the MECI geometry are shown in Figures 5d to 5g. Due to the asymmetric structure, the spectra show distinct XPS signatures for the C and N atoms of the NPhD and NPhA groups of the molecule. In the closed-shell state, the lowest-energy C and N K-edge ionization occurs in the PhD fragment, which exhibits significantly higher electron density and more efficient core-hole screening. Ionizing the PhA group requires 1 eV of additional ionization energy, significantly broadening the C and N K-edge XPS spectra in comparison to those of ground-state azobenzene (Figure 3). The XPS signatures of open-shell state more closely resemble the ground-state spectra with overlapping peaks originating from PhD and PhA. These results are consistent with the analysis of Mulliken charges (Figures 5b and 5c) and CASSCF natural orbitals (Figures S10 and S11), which reveal a more even electron density distribution between NPhD and NPhA in the state.
@@ 6 Conclusions
In this work, we presented an efficient implementation of multireference algebraic diagrammatic construction theory with core-valence separation for simulating core-ionized states and X-ray photoelectron spectra (CVS-IP-MR-ADC). Developed in the open-source and freely available Prism program,99 CVS-IP-MR-ADC takes advantage of spin adaptation, automatic code generation, and efficient handing of two-electron integrals via density fitting. Incorporating dynamic and static correlation, CVS-IP-MR-ADC allows to accurately simulate the X-ray or extreme ultraviolet photoelectron spectra (XPS) for molecules in excited electronic states or at nonequilibrium ground-state geometries. The CVS-IP-MR-ADC is resilient to intruder-state problems and allows to calculate many (10’s or even 100’s) excited states starting with a single CASSCF wavefunction for the reference electronic state.
We demonstrated the capabilities of our efficient CVS-IP-MR-ADC implementation by applying it to substituted ferrocene complexes and azobenzene molecule along its photoisomerization pathway. In all calculations, we used core-polarized quadruple-zeta basis sets, correlating all electrons in more than 1500 molecular orbitals. For the ground electronic states of substituted ferrocenes and trans-azobenzene, the carbon K-edge XPS spectra simulated using the extended second-order CVS-IP-MR-ADC method (CVS-IP-MR-ADC(2)-X) are in a very good agreement with experimental measurements. Encouraged by these results, we also computed the carbon and nitrogen K-edge XPS spectra for the azobenzene molecule in its excited states and at the geometry of asymmetric minimum energy conical intersection, which may be useful in interpreting the future time-resolved XPS experiments.
The efficient implementation strategy presented in this work can be used to develop fast MR-ADC methods for simulating other spectroscopic properties, including electron attachment and ionization in the UV/Vis region,37, 38 and neutral excitations in the UV/Vis and X-ray absorption spectroscopies (XAS).24, 35, 36 In addition, spin–orbit coupling effects can be incorporated to enable accurate simulations of L- and M-edge XPS and XAS spectra.203 These developments will be reported in the forthcoming publications of this series and will further expand the application domain of MR-ADC.
This work was supported by the National Science Foundation, under Grant No. CHE-2044648. The authors would like to thank Nicholas Gaba for testing the CVS-IP-MR-ADC code and valuable suggestions on its improvements. Computations were performed at the Ohio Supercomputer Center under Project No. PAS1583.225